Find the number of trailing zeros in M where M=201!-200!? answers 49,50,51,52
Answer by fcabanski(1391)

51


201! - 200! = 201*200! - 200! = 200*200!


To find the number of trailing zeros in any n! determine the number of times it is divisible by all smaller powers of 5.


The reason it works is that zeros happen whenever 10 is a factor, and 5 and 2 are factors of 10. Find the number of 5's that factor into the n! n, also the number of 5x5's and 5x5x5's etc.


Ignore remainders


200/5 = 40 + 200/25=8 + 200/125=1 = 49.


But this n! is multiplied by another number. Each of the zeros at the end of 200 will create a trailing 0. So it's 49+2 = 51.


If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com